# An adaptive finite element multigrid solver using GPU acceleration

### v2

@@ 1 Introduction
The combination of adaptive finite elements and geometric multigrid solvers is one of the most efficient approaches for the numerical approximation of partial differential equations. Adaptive mesh refinement schemes based on numerical a posteriori error estimates allow for optimal refinement [38] while geometric multigrid solvers provide the solution to the arising linear systems of equations in linear complexity. For many partial differential equations, the combination yields an overall approach with optimal computational complexity for a prescribed accuracy.
@@
The mathematical and computational optimality come at the price of complex algorithms and a lack of regular structure in the computations. This is due to the adaptivity of the discretization, resulting in an irregular mesh and stencils, and the multi-scale nature of the solvers. Furthermore, memory is often addressed in an unstructured way, hampering the parallelization of the computations.
@@
These challenges are one of the central reasons that parallel accelerator co-processors, such as GPUs and TPUs, are still only rarely used in numerical simulations, including for adaptive multi-grid methods. Accelerators are highly attractive since the computational power of CPUs has stagnated in the last decade while those accelerators has increased dramatically.
@@
Accelerators use task- and data-parallelism as well as specialization to achieve a very large peak performance. While a high degree of parallelism is inherent in some problems, e.g. in computer graphics and deep learning, it poses challenges for the implementation of adaptive algorithms in numerical linear algebra and scientific computing. One direction to address this gap is by changing to algorithms with highly regular computations, such as lattice Boltzmann methods in fluid mechanics [30, 32]. A minimally invasive use of GPUs in classical finite element simulation tools is also possible but yields little to no acceleration [16].
Another challenge for the the use of accelerator hardware in scientific computing is that these are optimized for single, half or even less precision while double precision remains the standard for most calculations in numerical mathematics. In [34], this is addressed by using a hierarchical approach to modify the linear systems, denoted as pre-handling. An alternative direction is the use of machine learning methods, especially deep neural networks, in the context of simulations, which can be executed very efficiently and in lower precision on accelerator hardware with speedups up to several orders of magnitude, e.g. [7].
For algebraic multi-grid methods, GPU accelerated libraries exist, e.g. amgx [27] and ginkgo [1]. However, these are likely less efficient than geometric multi-grid methods, if applicable.
@@
The integration of GPUs into an existing geometric multigrid solver was already presented in [17]. The obtained speedup was moderate but the work describes how a transparent implementation into an existing software can be achieved. The consequences of mixed precision arithmetic were also discussed in this work. In [15], the same authors presented a geometric multigrid method that is based entirely on sparse matrix-vector multiplications and can thus be easily and efficiently implemented on different hardware using suitable libraries. Applied to stationary linear differential equations, the solver exhibited up to a fivefold speedup relative to CPU-based systems.
In [19], the authors investigate the GPU parallelization of several smoothers within a matrix-based geometric multigrid method. The implementation allows for locally refined unstructured meshes, but, the obtained speedup is limited and does not substantially exceed that of multi-core CPU’s. NVIDIA [37] has presented a highly efficient geometric multigrid solver for 3d linear elliptic problems. It reaches very high throughput and demonstrates the potential of native CUDA implementations. It is, however, limited in terms of flexibility and its applicability to unstructured meshes.
@@
In this article, we show that highly efficient adaptive finite elements with local multigrid solvers can be implemented efficiently on GPU accelerators without limiting the flexibility of the method. For this, we formulate computations as matrix-vector and matrix-matrix computations. This allows us to use the cuBLAS and cuSPARSE libraries [29] that are highly optimized while the code remains close to those with BLAS calls on the CPU. Hence, the CPU and GPU versions can retain the same structure and differ only in the linear algebra function calls and additional CPU-to-GPU memory transfer in the GPU version. It, nonetheless, allows for the use of custom CUDA kernels when necessary and we demonstrate that the use of simple native CUDA code can provide significant speedups for operations that do not naturally map to (sparse) linear algebra. Due to these optimizations, the largest performance bottleneck comes from the memory transfer between CPU and GPU. In our final code, large parts of the computations have been transferred to the GPU and thus the slowdown through the transfer is also small.
@@
We demonstrate the generality and flexibility of our approach by applying it to two linear elliptic problems, namely the transport-diffusion equation in 2d and a 3d linear elasticity problem. The discretization of these two equations results directly in a linear system of equations that can be approximated with the GPU-accelerated geometric multigrid solver. As a third example, we consider the instationary Navier-Stokes equations. For these, we first derive an explicit pressure-correction method, which can be represented by matrix-vector multiplications and a pressure-Poisson problem, the latter one being solved with the multigrid solver. Special attention is given to the nonlinearity of the Navier-Stokes equations. Through a reformulation, also this will be approximated by a product with a pre-computed sparse matrix on the GPU. For all three examples, we obtain significant speed-ups between 5X and 20X for our final GPU parallelizations. At the same time, the systematic use of (sparse) linear algebra leads to easily understandable and maintainable code for the GPU computations.
@@
The remainder of the article is structured as follows. In Sec. 2 we introduce the mathematical notation and briefly present the finite element discretization. There, we also describe the geometric multigrid method. In Section 3, we will discuss accelerators such as GPUs and describe the specifics of the hardware and how they need to be reflected in the algorithms. We also introduce software libraries that facilitate the use of accelerator hardware. Section 4 describes our implementation of the multigrid process based on the cuSPARSE library [29]. Finally, in Section 5, we present numerical test problems and discuss the results. We conclude in Section 6.
@@ 2.1 Finite element discretization
[…]
@@
We assume shape regularity in the sense that $\|\nabla T_{T}\|\cdot\|\nabla T_{T}^{-1}\|\leq c$ uniform in $h>0$, see [33, Sec. 4.2]. Structural regularity is relaxed by allowing local mesh refinement with at most 1 hanging node per face, again, see [33, Sec 4.2] for the specific realization in the finite element library Gascoigne 3d [10].
Systems of partial differential equations such as the Navier-Stokes equations or elasticity will be discretized with equal-order finite elements, i.e. $\mathbf{V}_{h}=[V_{h}^{(r)}]^{c}$, where $c\in\mathbb{N}$ is the number of components. Gascoigne 3d combines these components locally. Taking the 3d Navier-Stokes equations as example this means that the vector is represented as a matrix with entries $x_{i,c}$ where $i$ refers to the grid node and $c$ to the component ($c=1$ pressure, $c=2,3,4$ velocities). The system matrix also has double indexing. Each matrix entry $A_{ij}$ is itself a (dense) matrix $A_{ij}\in\mathbb{R}^{n_{c}\times n_{c}}$ with $n_{c}=4$ in the case of 3d Navier-Stokes. Likewise, each vector entry $v_{i}$ is itself a vector $v_{i}\in\mathbb{R}^{n_{c}}$. Keeping the solution components together is beneficial in terms of cache efficiency for problems where one obtains large system matrices like Navier-Stokes ($n_{c}=4$) or 3d elasticity, where $n_{c}=6$ when the velocity and deformation fields are combined [8, 23]. The memory layout of this approach is not easily transferred to GPUs as standard libraries like cuSPARSE do not support such blocked matrices. Section 4.1.1 will give details.
@@ 2.2 Adaptivity and geometric multigrid hierarchy
[…]
@@
Step 5: Post-smooth return $\text{return }{\cal S}_{l}(x_{l}^{\prime\prime},b_{l})$
Adaptivity of the computations is realized by means of hierarchical local mesh refinement. If an element $T\in\Omega_{h}$ is chosen for refinement, it is replaced by 4 elements in 2d and 8 elements in 3d. We allow a level jump of 1 for neighboring elements. The actual refinement can therefore extend further into the domain. Faces and edges on elements with a level jump therefore have inner nodes that are unknowns only on one side. These nodes are called hanging nodes and are replaced by interpolations of their direct neighbours in the finite element approach. In the original Gascoigne, hanging node data used for this interpolation is stored in a sparse structure only covering exactly the hanging nodes. As preparation for GPU parallelization, the data is now represented in a sparse matrix. This matrix $\mathbf{H}_{h}$ is the diagonal identity in all rows belonging to standard degrees of freedom, whereas in rows belonging to hanging nodes, the matrix contains the interpolation weights such that the matrix vector product $\mathbf{y}=\mathbf{H}_{h}\mathbf{x}$ performs all interpolations at once. The computational effort is $O(N)$ ($N$ being the number of unknowns), regardless if hanging nodes are present. However, the constant is small and the overhead in relation to a treatment of the hanging nodes only is negligible. The transpose $\mathbf{H}_{h}^{T}$ plays an important role in the assembly of the matrix and variational residuals. If, for example, an integral $f_{i}=(f,\phi_{i})$ is calculated for $i=1,\dots,N$, all test functions $\phi_{i}$, including those in hanging nodes, are processed first. The result vector $\mathbf{f}_{h}=(f_{i})_{i}$ is then multiplied by $\mathbf{H}_{h}^{T}$ so that the test functions belonging to hanging nodes are correctly taken into account.
Mesh refinement can be based either on a priori knowledge or on a posteriori error estimates, the latter usually in the context of the dual weighted residual method [6, 8].
@@
Successive mesh refinement results in a hierarchy of meshes. Instead of using this hierarchy for the setup of the geometric multigrid solver, we start with the finest mesh $\Omega_{h}=:\Omega_{L}$ and recursively coarsen it until we reach a coarse mesh $\Omega_{H}=:\Omega_{0}$ with small complexity $N_{H}$. The mesh $\Omega_{H}$ is either the coarse starting mesh or any finer mesh chosen to give an optimal balance between recursive multigrid performance and fast direct solution of a small problem. By $L$ we denote the resulting number of multigrid layers. In each coarsening step, as many refinements as possible are taken back, yielding a hierarchy $\Omega_{h}=\Omega_{L}\succeq\Omega_{L-1}\succeq\dots\Omega_{0}=\Omega_{H}$. We follow a global multigrid approach where each intermediate level $\Omega_{l}$ spans the complete domain, see [4]. The main advantage of the global multi-grid is its simplicity. An approximation of the entire problem can be created on each grid level. This allows, for example, the introduction of global constraints, e.g. a normalisation of the pressure, $\int\Omega p\,\text{d}x=0$, in the context of incompressible flows or compliance with material balances $\sum_{i}c_{i}=1$ for chemical reactions, see [5]. The disadvantage of the global approach is the greater effort required, as the multigrid smoother covers the entire grid on each level. However, if one analyses the effort in terms of the accuracy that can be achieved with adaptive grids and not in terms of the number of degrees of freedom, it becomes clear that the global approach has no significant disadvantage in practical applications [4, 9, 8].
@@
The multigrid solver is the standard $V$-cycle. Coarse mesh problems are either solved exactly or approximated using a couple of smoothing steps. The geometric multigrid iteration is either used as linear solver or as pre-conditioner in a GMRES solver [35]. The GMRES solver is more robust and usually required for problems that are transport dominated or when meshes with large element aspect ratios are used.

### v3

@@ 1 Introduction
The combination of adaptive finite elements and geometric multigrid solvers is one of the most efficient approaches for the numerical approximation of partial differential equations. Adaptive mesh refinement schemes based on numerical a posteriori error estimates allow for optimal refinement [Stevenson2005] while geometric multigrid solvers provide the solution to the arising linear systems of equations in linear complexity [Hackbusch1985]. For many partial differential equations, the combination yields an overall approach with optimal computational complexity for a prescribed accuracy.
@@
The mathematical and computational optimality come at the price of complex algorithms and a lack of regular structure in the computations. This is due to the adaptivity of the discretization, resulting in an irregular mesh and stencils, and the multi-scale nature of the solvers. Furthermore, memory is often accessed in an unstructured way, hampering the parallelization of the computations.
@@
These challenges are one of the central reasons that parallel accelerator co-processors, such as GPUs and TPUs, are still only rarely used in numerical simulations, including adaptive multigrid methods. Accelerators are highly attractive since the computational power of CPUs has stagnated in the last decade while those accelerators has increased dramatically.
@@
Accelerators use task- and data-parallelism as well as specialization to achieve a very large peak performance. While a high degree of parallelism is inherent in some problems, e.g. in computer graphics and deep learning, it poses challenges for the implementation of adaptive algorithms in numerical linear algebra and scientific computing. One direction to address this gap is by changing to algorithms with highly regular computations, such as lattice Boltzmann methods in fluid mechanics [OBRECHT2013252, Ribbrock2010].
The performance characteristics of modern accelerator hardware, in particular GPUs, fundamentally differ from those of traditional CPUs. While GPUs offer massive parallelism and high arithmetic throughput, the performance in computations is in practice is often limited by bandwidth and latency between GPU-RAM and computational units, as well as by the cost of data transfers between host and device. These constraints pose significant challenges for classical finite element methods, which typically rely on sparse matrix storage and irregular memory access patterns to be efficient. As a consequence, a substantial body of recent work has focused on algorithmic reformulations of finite element methods that reduce memory traffic and increase arithmetic intensity to make more efficient use of GPUs. A key development in this direction is the use of matrix-free methods, where discrete operators are applied without explicitly assembling and storing sparse matrices. Instead, operator entries are recomputed on the fly, typically using tensor-product structure and sum-factorization techniques for higher-order elements. An overview of such approaches is given in [10.5555/3108096.3108097], where the potential of tensor-product evaluations and sum factorization for high-performance finite element computations is discussed, albeit not yet in the context of multigrid solvers. These ideas have since become central to GPU-oriented finite element design due to their favorable compute-to-memory ratio [Munch2023].
On GPUs, particularly high performance for multigrid methods has been demonstrated for fixed-stencil discretizations, which can be interpreted as a limit case of matrix-free computation. Thereby the application of an operator reduces to a small number of regular, local stencil operations with predictable memory access patterns. Recent studies such as [Antepara2024, Richardson2025] report excellent GPU efficiency for such methods, including for example $p$-multigrid techniques. While these results represent an important performance benchmark, their applicability to general, adaptive finite element methods is limited by the strong reliance on regular grid structures and fixed discretizations.
Also general finite element frameworks have made significant progress in leveraging GPUs. The MFEM library [Andrej2024] represents a modern, GPU-centric approach that supports matrix-free higher-order finite elements and provides $p$-multigrid preconditioning for Krylov subspace solvers. A distinguishing feature of MFEM is its partial realization of an end-to-end GPU workflow, in which most stages of the solver—including operator application and multigrid components—can be executed on the device, thereby minimizing data transfers. A similar direction is pursued in the deal.II library [Munch2023]. There GPU support has been steadily expanded within a matrix-free high-order finite element framework. Recent developments emphasize performance portability and the integration of accelerator backends into existing multigrid and matrix-free infrastructures. While these efforts demonstrate the feasibility of combining geometric multigrid with GPUs in general-purpose finite element libraries, they also highlight the complexity of designing algorithms that remain efficient across different hardware architectures and use cases.
A central difficulty in matrix-free multigrid methods on GPUs lies in the design of effective smoothers. Due to the lack of an explicitly assembled matrix, smoothers are often restricted to point-wise operations such as Jacobi or Chebyshev iterations. This may lead to reduced multigrid efficiency, especially for high-order discretizations [Munch2023]. Notable progress is in recent contributions on multigrid methods for high-order discontinuous Galerkin discretizations [Cui2025, Cui2025a], where efficient GPU-based vertex-patch smoothers are developed. These approaches demonstrate that carefully designed local solvers exploiting tensor-product structure can overcome some of the traditional limitations of matrix-free smoothing. The same concepts have also been successfully applied to the Stokes equations, using multigrid directly in the coupled velocity–pressure [Cui2025b]. In [Heuveline2012], the authors investigate the GPU parallelization of several smoothers within a matrix-based geometric multigrid method. The implementation allows for locally refined unstructured meshes, but, the obtained speedup is limited and does not substantially exceed that of multi-core CPUs
In contrast to matrix-free methods, matrix-based approaches on GPUs are less common, primarily due to the high memory footprint of sparse matrices and the irregular memory access patterns of sparse matrix–vector products. Nevertheless, these approaches remain attractive because they can often be integrated more easily into existing simulation frameworks. In [Xu2025], a two-level multigrid scheme with ILU-based smoothing on GPUs is investigated, demonstrating that meaningful acceleration can still be achieved for certain problem classes. A broader perspective on GPU usage in large-scale solvers is provided in the context of the PETSc framework [Mills2025]. The authors discuss solver design choices and scaling behavior toward exascale systems.
The primary advantage of matrix-free methods lies in their ability to drastically reduce memory usage by recomputing operator entries when needed. In situations where matrix-free solvers are not readily available—such as in complex multiphysics applications or legacy codes—mixed-precision techniques offer an alternative strategy. A general discussion of mixed-precision algorithms on GPUs, CPUs, and hybrid systems is given in [Abdelfattah2021]. In [RudaTurek] mixed-precision is made possible by using a hierarchical approach to modify the linear systems, denoted as pre-handling. In the context of multigrid solvers, mixed precision has been explored particularly for algebraic multigrid methods in [Tsai2023]. The potential of these approaches stems not only from reduced memory bandwidth requirements, but also from the significantly higher memory and compute throughput of low-precision arithmetic on modern GPUs.
For algebraic multigrid methods, GPU accelerated libraries exist, e.g. AmgX [Naumov2015] and Ginkgo [ginkgo-toms-2022]. NVIDIA [NVIDIAMG] has also presented a highly efficient geometric multigrid solver for 3d linear elliptic problems. A highly optimized pressure Poisson solver in GPUs is discussed in [GMRESAMG2024].
An alternative memory-reduction strategy orthogonal to matrix-free methods is proposed in [Boukhris2023]. There sparse matrices are represented in a block-wise manner using embedded stencil-like structures. This approach aims to combine some of the flexibility of matrix-based methods with the reduced memory footprint and regular access patterns typically associated with stencil computations.
Finally, a fundamentally different direction is the use of machine learning techniques, in particular deep neural networks, as surrogate models or solver components in scientific simulations. Such models can be executed very efficiently on accelerator hardware and can typically exploit from reduced numerical precision, leading to speedups of several orders of magnitude in specific applications, as demonstrated, for example, in [Bi:2023aa, Margenberg2024]. While these approaches are conceptually distinct from classical multigrid methods, they are increasingly explored as complementary tools in large-scale simulation workflows.
@@
A minimally invasive use of GPUs in classical finite element simulation tools is also possible but often yields little to no acceleration [5191718, Goddeke]. The obtained speedup was moderate but the work describes how a transparent implementation into an existing software can be achieved. The consequences of mixed precision arithmetic were also discussed in this work. In [GEVELER2013327], the same authors presented a geometric multigrid method that is based entirely on sparse matrix-vector multiplications and can thus be easily and efficiently implemented on different hardware using suitable libraries. Applied to stationary linear differential equations, the solver exhibited up to a fivefold speedup relative to CPU-based systems.
@@
In this article, we show that highly efficient adaptive finite elements with multigrid solvers can be implemented efficiently on GPU accelerators without limiting the flexibility of the method and without the large implementation effort that GPU-implementations can involve. For this, we use a matrix-based formulation that allows us to use the cuBLAS and cuSPARSE libraries [cuSPARSE] that are highly optimized while the code remains close to those with BLAS calls on the CPU. Hence, the CPU and GPU versions can retain the same structure and differ only in the linear algebra function calls and additional CPU-to-GPU memory transfer in the GPU version. It, nonetheless, allows for the use of custom CUDA kernels when necessary and we demonstrate that the use of simple native CUDA code can provide significant speedups for operations that do not naturally map to (sparse) linear algebra. Due to these optimizations, the largest performance bottleneck comes from the memory transfer between CPU and GPU, in particular, if system matrices are assembled on the CPU and then transferred to the GPU. In Section 5 we study numerical examples, where the repeated assembly can be avoided such that matrices are copied only once for all time steps. This reduces the overhead and the remaining time for memory transfer is negligible.
In our final code, large parts of the computations have been transferred to the GPU and thus the slowdown through the transfer is also small.
@@
We demonstrate the generality and flexibility of our approach by applying it to two linear elliptic problems, namely the transport-diffusion equation in 2d and a 3d linear elasticity problem. The discretization of these two equations results directly in a linear system of equations that can be approximated with the GPU-accelerated geometric multigrid solver. As a third example, we consider the instationary Navier-Stokes equations. For these, we first derive an explicit pressure-correction method, which can be represented by matrix-vector multiplications and a pressure-Poisson problem, the latter one being solved with the multigrid solver. Special attention is given to the nonlinearity of the Navier-Stokes equations. Through a reformulation, also this will be approximated by a matrix-vector multiplication with a pre-computed sparse matrix on the GPU. For all three examples, we obtain significant speed-ups between 5X and 20X for our final GPU parallelizations. At the same time, the systematic use of (sparse) linear algebra leads to easily understandable and maintainable code for the GPU computations.
@@
The remainder of the article is structured as follows. In Sec. 2 we introduce the mathematical notation and briefly present the finite element discretization. There, we also describe the geometric multigrid method. In Section 3, we will discuss accelerators such as GPUs and describe the specifics of the hardware and how they need to be reflected in the algorithms. We also introduce software libraries that facilitate the use of accelerator hardware. Section 4 describes our implementation of the multigrid process based on the cuSPARSE library [cuSPARSE]. Finally, in Section 5, we present numerical test problems and discuss the results. We conclude in Section 6.
Main contributions
We describe the integration of GPU acceleration into the Gascoigne 3d general-purpose finite element library. As the software did not consider accelerator hardware in its design, the focus is on making the most efficient use of the hardware while minimising interventions into the existing implementation. For this reason, we base our GPU acceleation on cuSPARSE and require custom kernels only for a small set of operations that cannot be expressed efficiently using linear algebra. Additionally, we present interpolation-based finite element techniques that enable us to formulate certain nonlinear terms with matrix-vector products based on sparse matrices with small stencils. This avoids numerical quadrature and extends the applicability of the cuSPARSE-based implementation to a larger class of problems. Our insight for the GPU acceleration is not specific to Gascoigne 3d and will hence likely generalize to other finite element libraries.
@@ 2.1 Finite element discretization
$Q^{r}:=\{x_{1}^{\alpha_{1}}\cdots x_{d}^{\alpha_{d}},\;\alpha_{i}\in\mathbb{N},\;0\leq\alpha_{i}\leq r\}$
@@
We assume shape regularity in the sense that $\|\nabla T_{T}\|\cdot\|\nabla T_{T}^{-1}\|\leq c$ uniform in $h>0$, see [Richter2017, Sec. 4.2]. Structural regularity is relaxed by allowing local mesh refinement with at most 1 hanging node per face, again, see [Richter2017, Sec 4.2] for the specific realization in the finite element library Gascoigne 3d [Gascoigne].
Systems of partial differential equations such as the Navier-Stokes equations or elasticity will be discretized with equal-order finite elements, i.e. $\mathbf{V}_{h}=[V_{h}^{(r)}]^{c}$, where $c\in\mathbb{N}$ is the number of components. Gascoigne 3d combines these components locally. Taking the 3d Navier-Stokes equations as example this means that the vector is represented as a matrix with entries $x_{i,c}$ where $i$ refers to the grid node and $c$ to the component ($c=1$ pressure, $c=2,3,4$ velocities). The system matrix also has double indexing. Each matrix entry $A_{ij}$ is itself a (dense) matrix $A_{ij}\in\mathbb{R}^{n_{c}\times n_{c}}$ with $n_{c}=4$ in the case of 3d Navier-Stokes. Likewise, each vector entry $v_{i}$ is itself a vector $v_{i}\in\mathbb{R}^{n_{c}}$. Keeping the solution components together is beneficial in terms of cache efficiency for problems where one obtains large system matrices like Navier-Stokes ($n_{c}=4$) or 3d elasticity, where $n_{c}=6$ when the velocity and deformation fields are combined [BraackRichter2006d, P1]. The memory layout of this approach is not easily transferred to GPUs as standard libraries like cuSPARSE do not support such blocked matrices. Section 4.1.1 will give details.
@@ 2.2 Adaptivity and geometric multigrid hierarchy
Step 0: Coarse mesh problem $\text{if }l=0$ $\text{ return }A_{0}^{-1}b_{0}$
@@
Step 5: Post-smooth $\text{return }{\cal S}_{l}(x_{l}^{\prime\prime},b_{l})$
Adaptivity of the computations is realized by means of hierarchical local mesh refinement. If an element $T\in\Omega_{h}$ is chosen for refinement, it is replaced by 4 elements in 2d and 8 elements in 3d. We allow a level jump of 1 for neighboring elements. The actual refinement can therefore extend further into the domain. Faces and edges on elements with a level jump therefore have inner nodes that are unknowns only on one side. These nodes are called hanging nodes and are replaced by interpolations of their direct neighbours in the finite element approach. In the original Gascoigne, hanging node data used for this interpolation is stored in a sparse structure only covering exactly the hanging nodes. As preparation for GPU parallelization, the data is now represented in a sparse matrix. This matrix $\mathbf{H}_{h}$ is the diagonal identity in all rows belonging to standard degrees of freedom, whereas in rows belonging to hanging nodes, the matrix contains the interpolation weights such that the matrix vector product $\mathbf{y}=\mathbf{H}_{h}\mathbf{x}$ performs all interpolations at once. The computational effort is $O(N)$ (NN being the number of unknowns), regardless if hanging nodes are present. However, the constant is small and the overhead in relation to a treatment of only the hanging nodes is negligible. The transpose $\mathbf{H}_{h}^{T}$ plays an important role in the assembly of the matrix and variational residuals. If, for example, an integral $f_{i}=(f,\phi_{i})$ is calculated for $i=1,\dots,N$, all test functions $\phi_{i}$, including those in hanging nodes, are processed first. The result vector $\mathbf{f}_{h}=(f_{i})_{i}$ is then multiplied by $\mathbf{H}_{h}^{T}$ so that the test functions belonging to hanging nodes are correctly taken into account.
Mesh refinement can be based either on a priori knowledge or on a posteriori error estimates, the latter usually in the context of the dual weighted residual method [BeckerRannacher2001, BraackRichter2006d].
@@
Successive mesh refinement results in a hierarchy of meshes. Instead of using this hierarchy for the setup of the geometric multigrid solver, we start with the finest mesh $\Omega_{h}=:\Omega_{L}$ and recursively coarsen it until we reach a coarse mesh $\Omega_{H}=:\Omega_{0}$ with small complexity $N_{H}$. The mesh $\Omega_{H}$ is either the coarse starting mesh or any finer mesh chosen to give an optimal balance between recursive multigrid performance and fast direct solution of a small problem. By $L$ we denote the resulting number of multigrid layers. In each coarsening step, as many refinements as possible are taken back, yielding a hierarchy $\Omega_{h}=\Omega_{L}\succeq\Omega_{L-1}\succeq\dots\Omega_{0}=\Omega_{H}$. To be precise, an element is coarsened, if it belongs to a group of four (in 3d eight) elements on the same mesh level that all arise from splitting the same common father element. Details are given in [BeckerBraack2000a]. We follow a global multigrid approach where each intermediate level $\Omega_{l}$ spans the complete domain, see [BeckerBraack2000a]. The main advantage of the global multigrid is its simplicity. An approximation of the entire problem can be created on each grid level. This allows, for example, the introduction of global constraints, e.g. a normalisation of the pressure, $\int_{\Omega}p\,\text{d}x=0$, in the context of incompressible flows or compliance with material balances $\sum_{i}c_{i}=1$ for chemical reactions, see [P2]. The disadvantage of the global approach is the greater effort required, as the multigrid smoother covers the entire grid on each level. However, if one analyses the effort in terms of the accuracy that can be achieved with adaptive grids and not in terms of the number of degrees of freedom, it becomes clear that the global approach has no significant disadvantage in practical applications [BeckerBraack2000a, J1, BraackRichter2006d].
@@
The multigrid solver is the standard $V$-cycle. Coarse mesh problems are either solved exactly or approximated using a couple of smoothing steps. The geometric multigrid iteration is either used as linear solver or as pre-conditioner in a GMRES solver [Saad1996]. The GMRES solver is more robust and usually required for problems that are transport dominated or when meshes with large element aspect ratios are used.